Find the critical angle for light traveling from a ruby (n=1.776) into air
1
Expert's answer
2020-03-25T11:01:28-0400

Critical angle from ruby in air


"sin(\alpha)=\frac{1}{n}""\alpha=arcsin(\frac{1}{n})=arcsin(\frac{1}{1.776})=34^o"